The current in (or through) the 10 kΩ resistor is:

a.

0.5 mA

b.

2.0 mA

c.

2.4 mA

d.

10 mA

(Schematic, which shows a 24 volt source in series with a 10 kΩ resistor and a 2.0 kΩ resistor )
Answer #1

Using KCL: 24=I*10k+I*2k then

Current I=24/(10k+2k)= 24/12k = 2mA

So the answer is 2 milli amperes

Go for option b